Solve using elimination or substitution 4x-7y =-16;6x+3y = 30?

Answer by oberobic(2304)

4x -7y = -16
6x +3y = 30
.
Multiply the first equation by 3 and the second equation by 2, then subtract to eliminate the 'x' terms.
.
12x -21y = -48
12x +6y = 60
--------------
-27y = -108
y = 4
.
substitute y=4 into first equation
.
4x -7(4) = -16
4x = -16 +28
4x = 12
x = 3
.
Always check your work.
Substitute these values in the second equation.
.
6x +3y = 30
6(3) = 18
3(4) = 12
18+12 = 30
Correct.
.
Answer: x=3, y=4
.
Done.